OpenStudy (anonymous):

Combine: sqrt 75a^3 + a sqrt 12a

OpenStudy (anonymous):

= sqrt75*a^3/2 + a*sqrt12*sqrta = sqrt75*a^3/2 +a*a^1/2*sqrt12 =sqrt75*a^3/2 + sqrt12*a^3/2 =(sqrt75+sqrt12)a^3/2